# Unsupervised Text Embedding Space Generation Using Generative Adversarial Networks for Text Synthesis

### v1

@@ 5 Results
We use FBD, MSJ, LM, and DSR for comparing the results of the synthesized text at every epoch of adversarial training. Since FBD and MSJ require a real text set, the test set is used as the real text set. DSR is calculated with the training set as the data memorization ratio needs to be computed. FBDs of the TESGAN-based models are lower than MLE, while baselines increase after having the lowesIn terms of MSJ, as shown in Figure 1(b)-1(e), the previous studies report lower values than the TESGAN-based models at the end of adversarial training, despite having higher results in the beginning. On the other hand, MSJ results of the TESGAN-based models slightly increase during adversarial training. Moreover, some MSJ5 results of the original TESGAN are higher than MLE during adversarial training as shown in Figure 1(e). In the case of DSR shown in Figure 1(f), DSR results of the original TESGAN increase during adversarial training, and some results are higher than MLE. On the other hand, the results of the previous studies decrease during adversarial training, resulting in lower values than the TESGAN-based models in the end. As adversarial learning progresses, the results of the baselines deteriorate because the LSTM generator tends to generate only a few unique sentences.
We choose the best model of each method considering the overall results shown in Figure 2. As shown in Table 2, TESGAN and P-TESGAN at the end of 17 and 10-epoch of adversarial training are selected, respectively. For all the baselines, the 1-epoch models are selected immediately after supervised pre-training. After selecting the best models, we calculate the LM based on the text generated by each model, and the TESGAN-based models show the best results. The results of Gaussian random noise demonstrate that the TESGAN results are attributed to the seeds from the generator. Table 3 shows the synthetic text by TESGAN and Gaussian random noise.
In conclusion, MLE is a supervised pre-trained generator applied before adversarial training, but most of the result curves of the prior methods show lower performance than MLE during adversarial training. On the other hand, our TESGAN-based models have shown better results than MLE or have improved performance during adversarial training.
Error Analysis
@@
Figure 3: Illustration of TESGAN training results. TESGANs show similar trends for every trial and SEMs decrease during adversarial training.
We also conduct three additional TESGAN training without setting a manual seed in the code to confirm the reproducibility. As a result, the overall tendency of training outcomes during adversarial training is similar. We compute the mean and the SEM of the four results and the SEM is calculated by $\frac{\sigma}{\sqrt{n}}$ (n denotes the number of experiments). Also, the SEM of each epoch decreases during adversarial training, which means that each TESGAN converges. Figure 3 shows the results of four experiments, mean, and SEM.
@@ 6 Discussion
In this experiment, we find that the performance of TESGAN depends on hyperparameters and training strategies. In the generator, the ReLU and the hyperbolic tangent (tanh) are worse than the Leaky ReLU [19]. Furthermore, deep convolutional layers and batch normalization tend to synthesize monotonous text. Thus, layers of the generator are constructed not deeply but widely and do not have batch normalization. Further discussion on the generator is in Appendix A.
It is important to find appropriate steps of the models because the discriminator cannot follow the learning of the generator if the number of generator-step (g-step) is too large compared to the number of discriminator-step (d-step) and vice versa. Commonly, since the convergence of the generator is slower than that of the discriminator, the g-step and the d-step are set to two and one, respectively. The entire pseudocode and optimization study are explained in Appendix B.
Ablation Study
Method FBD $\downarrow$ MSJ2 $\uparrow$ MSJ3 $\uparrow$ MSJ4 $\uparrow$ MSJ5 $\uparrow$ DSR normal-$\uparrow$ LMnormal-$\ast$ $\downarrow$
@@
We experiment to see the effect of the four objective functions in Section 3.4 and the results are shown in Table 4. When SOD and SDP are not used, there is a big difference in the results, indicating that the SOD and SDP are important for high-quality text synthesis. Since MSJ evaluates text based on the n-grams of tokens, the order of the synthesized text is important. Accordingly, the MSJ results of the w/o SOD are worse than those of the w/o SSD, which proves that SOD can capture the token order representations. Lastly, the results are generally high when all four objective functions are used.
Activation Function Study
Activation FBD $\downarrow$ MSJ2 $\uparrow$ MSJ3 $\uparrow$ MSJ4 $\uparrow$ MSJ5 $\uparrow$ DSR normal-$\uparrow$ LMnormal-$\ast$ $\downarrow$
@@
The results are different depending on the activation functions used at the end of the seed making process. We conduct experiments on sigmoid, tanh, and non-use cases during the seed making process, and their results are shown in Table 5. Table 6 shows the quality of the synthetic text for tanh and non-use cases, and the results are worse than those using sigmoid in Table 3. However, according to Table 5, DSR results of non-use case is higher than the sigmoid case. Thus, we can see that a higher DSR does not always mean good quality because DSR only considers data memorization. Therefore, we select the model using the sigmoid activation function, which has better results for FBD and MSJ, and moderately high DSR.
Case Study
Figure 4: DSR results according to the number of seed interpretation model parameters.
Pre-trained GPT-2 which has 124M parameters used as the seed interpretation model is a powerful language model. Therefore, we need to confirm whether the low data memorization comes from transfer learning or the TESGAN framework. We trained three smaller seed interpretation models from scratch to measure the data memorization and they have 54M, 75M, and 96M parameters each. As shown in Figure 4, DSR is high regardless of the number of model parameters during adversarial training. The results in Figure 4 indicate that the low data memorization comes from the TESGAN framework.
@@ 7 Conclusion
Herein, we proposed a novel unsupervised text synthesis framework, TESGAN. TESGAN enabled the gradient backpropagation of natural language discrete tokens by creating a continuous text embedding space called a seed. In most text-GAN studies, data memorization had been inevitable because the generator have to be pre-trained with an autoregressive way before adversarial training. Therefore, we suggested TESGAN, which removed the data memorization issue by applying an unsupervised GAN framework that does not directly refer to the training data. TESGAN improved text synthesis performance during adversarial training and resulted in the best FBD and LM. Also, TESGAN showed the lowest data memorization ratio and the case study proved that these results are attributed to the TESGAN framework. Furethermore, ablation study shows the importance of the four objective functions. This paper implies the potential of continuous embedding spaces as well as discrete tokens for text synthesis with unsupervised learning. By combining the perspective of viewing text as a continuous space with publicly available Large Language Models [25], models will synthesize more expressive sentences, and we hope many follow-up studies will come out.
Appendix 0.A Details of Models
(a) Generator
(b) Seed Structure Discriminator (SSD)
(c) Seed Order Discriminator (SOD)
(d) Seed Distribution Prediction (SDP)
Figure 5: Illustrations of the generator, two discriminators, and SDP. In the generator, P-TESGAN makes perturbed seeds by adding zero-centered normal distribution noise $z$ (gray) to the output (blue) from the generator. SSD predicts whether the seed is real or fake using the $[CLS]$ special token. SOD considers both forward and backward contexts of the seed. SDP is used to improve the fake seeds of the generator during adversarial training by reducing the distance between real and fake seed distributions.
0.A.1 Seed Interpretation Model
@@
The seed interpretation model $f_{\theta}(\cdot)$ is necessary to predict subsequent sentences a seed makes. Therefore, the seed interpretation model must be trained with multi-turn sentences in an autoregressive way Our interpretation model inherits the 12-layer GPT-2, derived from the decoder of the transformer language model [26], and has 124M parameters. We used the model achieving the highest NLTK BLEU-4 in the validation set in Table 1 as the seed interpretation model.
0.A.2 Generator
@@
The generator $g_{\phi}(\cdot)$ consists of two 1D transposed convolutional layers and has 3.3M parameters. The first and the second layers conduct convolution with 128 and 16 filters, respectively. Since sentences vary according to types of tokens and their order, forms of the real seed $H_{real}$ are also varied. The generator generates seeds from the uniform distribution noise $X$ with an interval of $[-10,10)$ to make diverse forms of the seeds. Furthermore, the fake seeds, results of deep convolutional layers and batch normalization, tend to synthesize only monotonous sentences. Thus, layers of the generator are constructed not deeply but widely and the generator does not have batch normalization layers.
Leaky ReLU is used as the activation function between the two convolutional layers. When the ReLU is used, Dying ReLU occurs; the negative values become zero, making it unsuitable for text synthesis where diversity is important. Also, tanh is not adequate because of the gradient vanishing problem. Figure 4(a) shows convolutional processes of the generator.
0.A.3 Seed Structure Discriminator (SSD)
Sentence structure is important for constructing a complete sentence. For example, “I love you so much” is structurally error-free, but “I love like so much” and “I love” are not. Because real seeds are created from perfect sentences, they retain the structural representation of sentences. Therefore it is important that the fake seeds should capture the structural features of the real seeds.
@@ 0.A.3 Seed Structure Discriminator (SSD)
We assume that every sentence can be the first sentence in multi-turn cases. Thus, the real seeds are obtained from sentences where the $[CLS]$ token is inserted at the beginning. We use the 2-layer BERT $d_{\alpha}(\cdot)$ to capture the structural features of sentences, and the $[CLS]$ token is used to predict whether the seed is real (label 1) or fake (label 0). In addition, real and fake seeds do not pass through the embedding part of the BERT because they are already embedding spaces. Finally, the BERT used in SSD has 54M parameters and SSD is used to update both the discriminator and the generator during adversarial training. The detailed figure of SSD is shown in Figure 4(b).
0.A.4 Seed Order Discriminator (SOD)
@@
The order of tokens is important for constructing sentences. For example, “I love you so much” is syntactically correct, but “I you love so much” and “I love you much so” are not. To do this, we use a 2-layer Bidirectional LSTM to consider both forward and backward directions of sentences and the model has 24M parameters. Concatenated hidden states of the last token ($[SEP]$ or $[PAD]$) and the first token ($[CLS]$) are used to predict whether the seed is real (label 1) or fake (label 0). Finally, SOD updates both the discriminator and the generator during adversarial training. The detailed figure of SOD can be found in Figure 4(c).
Appendix 0.B Reproducibility
0.B.1 Optimization

### v4

@@ 5.1 Metric-based Evaluation
The FBDs of the TESGAN-based models are lower than the MLE-based autoregressive results, while the baselines increase after having the lowest value at the end of the first epoch of adversarial training, as shown in Figure 5(a). In terms of MSJ, as shown in Figure 5(b)-5(e), the previous studies report lower values than the TESGAN-based models at the end of adversarial training, despite having higher results in the beginning. On the other hand, MSJ results of the TESGAN-based models slightly increase during adversarial training. Moreover, some MSJ5 results of the original TESGAN are higher than MLE-L during adversarial training, as shown in Figure 5(e). In the case of DSR, as shown in Figure 5(f), the original TESGAN also increases during adversarial training, and some results are higher than MLE-L. On the other hand, the results of the previous studies decrease during adversarial training, resulting in lower values than the TESGAN-based models in the end. As adversarial learning progresses, the results of the baselines deteriorate because the LSTM generator tends to generate only a few unique sentences. Furthermore, we will explain the reasons why the MLE-G results of the GPT-2 base are relatively poor in the following section.
We chose the best model of each method considering the FBD, MSJ, and DSR results because these metrics evaluate quality, diversity, and data memorization. We evaluated the text generated by each model per epoch using the metrics mentioned above and compared the best-performing models [^5] Detailed results are shown in Appendix E. According to Table 8 in Appendix E, we compare the baselines at 1-epoch with TESGAN and P-TESGAN at 17 and 10-epochs, respectively. After selecting the best models, we calculated the Language Model score (LM) and Self-BLEU (SBL) based on the text generated by each model. As shown in Table 1, the TESGAN-based models show the highest results in FBD and DSR. Also, the TESGAN-based models show comparable results in MSJ compared to the baselines and display the highest results among the adversarial-based methods in terms of LM score. However, in terms of SBL, TESGAN-based models perform worse than the baselines. In addition, the results of Gaussian random noise demonstrate that the TESGAN results are attributed to the seeds from the generator. The first group of Table 2 shows the synthetic text by TESGAN and Gaussian random noise. In conclusion, MLE-L is a supervised pre-trained generator applied before adversarial training, but most of the result curves of the prior methods showed lower performance than MLE-L during adversarial training. On the other hand, our TESGAN-based models showed better results than MLE-L or improved performance during adversarial training. Finally, the results according to the epoch of the LM and SBL of each model are shown in Appendix E.
5.2 Analysis of Autoregressive Models
In this section, we will analyze the results of the MLE-based autoregressive models. Other baseline models pretrain an LSTM-based generator before starting adversarial training, while the TESGAN framework employs a GPT-2-based pre-trained seed interpretation model. The results of the MLE-based models in Section 5.1 are based on the evaluation of corpora generated in an autoregressive manner using the two pre-trained models. MLE-based models generate sentences in an autoregressive manner, starting from a specific $[start\_token]$ and predicting the next token. If the model predicts the next token in a greedy manner, all generated sentences would be identical, exhibiting deterministic behavior. To prevent this, MLE-based models sample the next token based on the probability logits Yu et al. (2017). This way, MLE-L results in diverse token choices since the logit probability differences are not large. On the other hand, MLE-G training fits the data better than LSTM-based models, resulting in significantly larger differences in the logits of the next token. As a consequence, MLE-G is relatively deterministic compared to MLE-L. Therefore, when generating sentences without the use of the softmax temperature technique Hinton et al. (2015), MLE-G delivers high quality, but it struggles to produce a variety of sentences. In practice, sentences generated by MLE-G lack diversity, which led to relatively lower results in Section 5.1. However, it is worth noting that while diversity may be lacking, the quality of the generated sentences is high, and this aspect will be demonstrated in the next section.
5.3 Human Evaluation
Method Avg. Score
TESGAN (ours) 4.2
P-TESGAN (ours) 3.4
SeqGAN Yu et al. (2017) 2.4
RankGAN Lin et al. (2017) 2.0
MaliGAN Che et al. (2017) 1.0
PG-BLEU Yu et al. (2017) 1.6
MLE-L Yu et al. (2017) 3.0
MLE-G 3.8
Table 3: Human evaluation scores (1 $\sim$ 5).
We conducted human evaluations based on the corpora generated by each model. The corpora, comprising 50 randomly selected unique sentences that do not duplicate those from the training set, were assessed by 10 annotators. We asked annotators to give higher scores to corpora that contained more natural sentences on a scale from 1 to 5. The scores presented in Table 3 represent the average scores assessed by each person. According to Table 3, TESGAN received the highest score, with MLE-G achieving the second-highest result. As mentioned in Section 5.2, MLE-G, despite facing challenges in generating diverse sentences, was able to produce high-quality sentences.
5.4 General Applicability
(a) Zero-shot results
(b) Non-zero-shot results
(c) DSR, LM, SBL results
Figure 7: Zero-shot, non-zero-shot results of IMDb-trained TESGAN. DSR, LM, and SBL results of IMDb-trained TESGAN are normalized for ease of viewing.
In this section, we trained TESGAN with IMDb using a larger volume than DailyDialog to assess the general applicability of TESGAN. The IMDb-trained TESGAN is evaluated with both the DailyDialog test set (zero-shot) and the IMDb test set (non-zero-shot). Figures 6(a) and 6(b) display the zero-shot and non-zero-shot test results of the IMDb-trained TESGAN, respectively. Additionally, the zero-shot results generally exhibit a similar trend to the non-zero-shot test, suggesting that the model is being trained without bias toward the training data. Furthermore, the second group in Table 2 presents the synthetic text results of the IMDb-trained TESGAN. Both the zero-shot and text synthesis results indicate that TESGAN’s outcomes do not vary significantly depending on the dataset, implying that TESGAN generalizes well and can be trained on diverse datasets. Figure 6(c) also illustrates the DSR, LM, and SBL results of the IMDb-trained TESGAN. Since these metrics are evaluated not on the test set but on generated text data, they consistently yield results regardless of the zero-shot test.
5.5 Error Analysis
@@
Figure 8: Illustration of TESGAN training results. TESGANs show similar trends for every trial, and SEMs decrease during adversarial training.
We also conducted three additional TESGAN trainings without setting a manual seed in the code to confirm reproducibility. To assess whether the sentences generated by the model for each trial converge as adversarial training progresses, we calculated the Standard Error of the Mean (SEM) based on the average results. SEM is equivalent to the standard deviation of a sample mean taken from a population and represents the standard deviation that indicates the extent of variability in sample means. SEM is calculated by $\frac{\sigma}{\sqrt{n}}$ ($\sigma$ and $n$ denote average results and the number of trials). As a result, the overall tendency of training outcomes during adversarial training is similar. Furthermore, the SEM of each epoch decreases during adversarial training, indicating that each TESGAN converges. Figure 8 displays the results of the four experiments, including the average results and SEM.
@@ 6 Discussion
6.1 Generator and Training Strategy
We found that the performance of the TESGAN framework depends on the generator’s architecture. When ReLU was used, dying ReLU Lu et al. (2019) occurred, where the negative values became zero, making it unsuitable for text synthesis where diversity is important. Additionally, the hyperbolic tangent (tanh) was not adequate due to the problem of gradient vanishing Wang et al. (2019). Consequently, we adopted Leaky ReLU Maas et al. (2013) as the activation function between two convolutional layers of the generator. Furthermore, deep structures and batch normalization tended to result in monotonous text synthesis. Therefore, we designed the generator’s layers to be wide rather than deep without batch normalization.
We also observed that the convergence of TESGAN depends on the parameter update rate of the discriminators and the generator. As in Algorithm 1, the discriminators update their parameters only during odd training epochs to allow the generator to catch up with the discriminator’s learning because the convergence of the generator is commonly slower than that of the discriminator. When the discriminators updated their parameters at every training epoch, the same as the generator, adversarial training became unbalanced. Additionally, we conducted further experiments by changing the update frequency of the generator from once to three times per mini-batch step. When the generator updated only once per step, the same as the discriminators, it could not keep up with the learning of the discriminators. On the other hand, when the generator updated three times per step, the discriminators could not keep up with the learning of the generator. Therefore, we chose to update the generator twice per step, resulting in the generator being updated four times more frequently per two epochs than the discriminators, as explained in Algorithm 1.
6.2 Ablation Study
[…]
@@ 6.2 Ablation Study
In this section, we confirm the effect of the four objective functions in Section 3.4, and the results are shown in Table 4. When Seed Order Discriminator (SOD) and Seed Distribution Prediction (SDP) were not used, there was a significant difference in the results, indicating that SOD and SDP are important for high-quality text synthesis. Since MSJ evaluates text based on the n-gram of tokens, the order of the synthesized text is important. Accordingly, the MSJ results of the ”w/o SOD” in Table 4 are worse than those of the ”w/o Seed Structure Discriminator (SSD)”, which proves that SOD can capture the token order representations. The four objective functions were used to achieve a good overall result, demonstrating that each of the four objective functions is playing a unique role.
6.3 Activation Function Study
[…]
@@ 6.3 Activation Function Study
The results varied depending on the activation functions used at the end of the seed-making process. We conducted experiments on sigmoid, tanh, and non-use cases during the seed-making process, and their results are shown in Table 5. Table 6 shows the quality of the synthetic text for tanh and non-use cases, and the results are worse than those using sigmoid in Table 2. However, according to Table 5, the DSR results of the non-use case are higher than the sigmoid case. Thus, we can see that a higher DSR does not always mean good quality because DSR only considers data memorization. Therefore, we select the model using the sigmoid activation function, which has better results for FBD and MSJ, and moderately high DSR.
6.4 Data Memorization Study
Figure 9: DSR results according to scales of seed interpretation model.
The pre-trained GPT-2, which has 124M parameters and is used as the seed interpretation model, has been trained on relatively large corpora. Therefore, we need to confirm whether the low data memorization comes from transfer learning or the TESGAN framework. We trained three smaller seed interpretation models from scratch to measure the data memorization and they have 54M, 75M, and 96M parameters each. As shown in Figure 9, DSR is high regardless of the number of model parameters during adversarial training, indicating that the low data memorization comes from the TESGAN framework.
@@ 7 Conclusion
In this work, we proposed a novel unsupervised text synthesis framework, TESGAN. TESGAN facilitated the gradient backpropagation of natural language discrete tokens by creating a continuous text embedding space called a seed. In most text-GAN studies, data memorization had been inevitable because the generator had to be pre-trained with an autoregressive approach before adversarial training. Therefore, we introduced TESGAN, which mitigated the data memorization issue by applying an unsupervised GAN framework that does not directly refer to the training data. TESGAN improved text synthesis performance during adversarial training and resulted in the best or comparable results in terms of evaluation metrics. Additionally, TESGAN exhibited the lowest data memorization ratio, and the data memorization study confirmed that these results were attributable to the TESGAN framework. Furthermore, TESGAN achieved the highest scores in human evaluations. The ablation study highlighted the importance of the four objective functions, and the synthetic text results from a large dataset-trained TESGAN demonstrated its general applicability. This paper underscores the potential of continuous embedding spaces in conjunction with discrete tokens for text synthesis through unsupervised learning. By integrating the concept of viewing text as a continuous space with publicly available Large Language Models Touvron et al. (2023), models can synthesize more expressive sentences, and we anticipate that many follow-up studies will emerge.
Appendix A Details of Models
A.1 Seed Interpretation Model
@@
The seed interpretation model $f_{\theta}(\cdot)$ is necessary to predict subsequent sentences a seed makes. Therefore, the seed interpretation model must be trained with multi-turn sentences in an autoregressive way. Our interpretation model inherits the 12-layer GPT-2, derived from the decoder of the transformer language model Vaswani et al. (2017), and has 124M parameters. We used the model achieving the highest NLTK BLEU-4666https://www.nltk.org/_modules/nltk/translate/bleu_score.html in the validation set of each dataset in Table 7 as the seed interpretation model.
A.2 Generator
@@
The generator $g_{\phi}(\cdot)$ consists of two 1D transposed convolutional layers and has 3.3M parameters. The first and the second layers conduct convolution with 128 and 16 filters, respectively. Since sentences vary according to types of tokens and their order, forms of the real seed $H_{real}$ are also varied. The generator generates seeds from the uniform distribution noise $X$ with an interval of $[-10,10)$ to make diverse forms of the seeds. Furthermore, the fake seeds generated by the deep convolutional layers and batch normalization results tend to synthesize only monotonous sentences. Thus, layers of the generator are constructed not deeply but widely and the generator does not have batch normalization layers. Leaky ReLU is used as the activation function between the two convolutional layers.
A.3 Seed Structure Discriminator (SSD)
@@
Sentence structure is important for constructing a complete sentence. For example, “I love you so much” is structurally error-free, but “I love like so much” and “I love” are not. Because real seeds are created from perfect sentences, they retain the structural representation of sentences. Therefore it is important that the fake seeds should capture the structural features of the real seeds. We assume that every sentence can be the first sentence in multi-turn cases. Thus, the real seeds are obtained from sentences where the $[CLS]$ token is inserted at the beginning like Equation 3. We use the 2-layer BERT $d_{\alpha}(\cdot)$ to capture the structural features of sentences, and the $[CLS]$ token’s feature is used to predict whether the seed is real (label 1) or fake (label 0). In addition, real and fake seeds do not pass through the embedding part of the BERT because they are already embedding spaces. Finally, the BERT used in SSD has 54M parameters.
A.4 Seed Order Discriminator (SOD)
@@
The order of tokens is important for constructing sentences. For example, “I love you so much” is syntactically correct, but “I you love so much” and “I love you much so” are not. We use a 2-layer Bidirectional LSTM to consider both forward and backward directions of sentences and the model has 24M parameters. The concatenated hidden states of the last token ($[SEP]$ or $[PAD]$) and the first token ($[CLS]$) are used to predict whether the seed is real (label 1) or fake (label 0).
Appendix B Loss Function
Here, we show whole loss functions of TESGAN:
Seed Interpretation $\textit{Seed Interpretation Model}:\\ \mathcal{L}_{LM}=-\frac{1}{N}\sum^{N}_{n=1}{\log{\frac{exp(x_{n,y_{n}})}{\sum^{C}_{c=1}{exp(x_{n,c})}}}}\\ \\ \textit{Adversarial Training Training}\\ \mathrm{d}-\mathrm{step}:\\ \mathcal{L}_{D}=-\frac{1}{N}\sum_{i=1}^{N}\Big{\{}\big{[}y_{i}\log{x_{i}}+(1-y_{i})\log{(1-x_{i})}\big{]}_{SSD}^{real/fake}+\\ \big{[}y_{i}\log{x_{i}}+(1-y_{i})\log{(1-x_{i})}\big{]}_{SOD}^{real/fake}\Big{\}}\\ \\ \mathrm{g}-\mathrm{step}:\\ \mathcal{L}_{G}=-\frac{1}{N}\sum_{i=1}^{N}\Big{\{}\big{[}y_{i}\log{x_{i}}+(1-y_{i})\log{(1-x_{i})}\big{]}_{SSD}^{fake}+\\ \big{[}y_{i}\log{x_{i}}+(1-y_{i})\log{(1-x_{i})}\big{]}_{SOD}^{fake}+\\ \sigma\big{(}f_{\theta}(H_{real})\big{)}\log{\frac{\sigma\big{(}f_{\theta}(H_{real})\big{)}}{\sigma\big{(}f_{\theta}(H_{fake})\big{)}}}+\\ \|\mu_{r}-\mu_{f}\|_{2}^{2}+\|H_{real}-H_{fake}\|_{1}\Big{\}}\$ (15)
The first loss function in Equation 15 is cross-entropy and is used to train the seed interpretation model. The loss functions used in adversarial training operate differently in the discriminator and generator steps. In the discriminator step (d-step), the loss function is designed to train the discriminator to distinguish between real and fake seeds, predicting them as 1 and 0, respectively. On the other hand, in the generator step (g-step), the loss function aims to train the generator to predict fake seeds as 1. Additionally, SDP and SFP losses are added to assist the generator learning during the g-step.
Appendix C Statistics of Datasets
Table 7 shows the statistics of the two datasets used in this paper. We excluded single-turn reviews when constructing the IMDb multi-turn dataset. For the baseline performance experiments, we generated fake seeds equal to the number of sentences in the DailyDialog dataset to conduct TESGAN training (adversarial training). Similarly, for the general applicability experiments, we generated nearly 300k fake seeds to conduct the experiments with IMDb datasets.
Statistics DailyDialog IMDb
Train Validation Test Train Validation Test
##\## of multi-turn set 11,118 1,000 1,000 24,890 12,500 12,390
Total sentences 87,170 8,069 7,740 299,137 150,369 148,768
Avg. ##\## of turns per set 7.84 8.07 7.74 12.02 12.03 12.01
Avg. ##\## of words per sentence 11.30 11.21 11.44 19.34 19.40 19.28
Avg. ##\## of tokens per sentence 14.51 14.39 14.69 24.25 24.31 24.20
Table 7: Statistics of datasets
Appendix D Hyperparameters
The TESGAN framework has two training steps. The first step is seed interpretation model training. The multi-turn data for seed interpretation model training were limited to a maximum of four and eight turns in performance (DailyDialog-trained) and general applicability (IMDb-trained) experiments, respectively. Also, the maximum length of the sentence was set to 16 and 32 for each experiment (total sequence length of each experiment was 64 and 256). The 12-layer GPT-2 is used as the seed interpretation model, and both hidden and embedding dimensions are 768. We adopted the byte-pair-encodings (BPE) Sennrich et al. (2016) tokenizer with 50,260 vocabularies in the seed interpretation model. We used the Adam optimizer Kingma and Ba (2014) with $1e^{-3}$ learning rate to train the seed interpretation model and set the mini-batch size to 100.